OpenStudy (anonymous):

-5/8÷6 divide & simplify Just want to double check my work..thanks

OpenStudy (anonymous):

\[\frac{-5}{8}\div 6=\frac{-5}{8}\times\frac{1}{6}=\frac{-5}{48}\]

OpenStudy (triciaal):

@mzfreda2u next time post your work to be checked otherwise you are just waiting for an answer. so, did you get the same as @Orion1213?